OpenStudy (anonymous):

I have one last question... Juan always saves the same amount from his weekly allowance. The table shows how much he has saved at different times. How much will Juan have saved in 14 wk? Juan’s Savings Weeks 3 6 9 12 Amount saved 19 28 37 46 I just want to check my answer and see if its correct

OpenStudy (anonymous):

I think the answer is $52

OpenStudy (anonymous):

@AloneS are you on?

OpenStudy (anonymous):

im going with 52 I guess...

alones (alones):

yeah i wasa in biology :/

OpenStudy (anonymous):

@Vocaloid hope im not asking to much for help :) I just want to understand this, I think its $52

OpenStudy (anonymous):

thanks @AloneS for the help too XD

OpenStudy (anonymous):

@Sanic I need help

alones (alones):

add 6$ to48 and that's you're answer :)

OpenStudy (anonymous):

54?

OpenStudy (anonymous):

or is it 52? 54 or 52...?

alones (alones):

okayi i wanna know where did you get 52??

OpenStudy (anonymous):

I mean, what should I go with?

OpenStudy (anonymous):

Y = ax + b the difference in numbers is 9 and the difference in weeks is 3 so a = 9/3 = 3 b = 19 - 9 = 10 the equation is y = 3x + 10 sub 14 into equation as x y = (14 x 3) + 10 y = 42 + 10 y = $52

alones (alones):

well yeah i know but maybe u right

OpenStudy (anonymous):

I'll go with that and tell you what I get...

OpenStudy (anonymous):

welp, I submited it....

alones (alones):

nov u should go with you'rs i'm kinda confused here

OpenStudy (anonymous):

I got an 80% I passed :)

OpenStudy (anonymous):

now as i always do, put the right answers for others that need help on this too XD

OpenStudy (anonymous):

Anne runs at a constant rate. The table shows the distance she traveled at specific times during her run. What is the slope of the line that represents Anne’s speed? Anne’s Speed Minutes 5 10 15 20 Distance (mi) 0.625 1.25 1.875 2.5 --->(0.125)

OpenStudy (anonymous):

Carmen hikes at a constant rate. The graph shows the distance Carmen has traveled at different times during his hike. https://static.k12.com/calms_media/media/1512000_1512500/1512347/1/eb3b3cff79e8e8ac2e4fdb05b3282a81286ddb6a/MS_IMC-140516-100011.jpg What is the slope of the line that represents the situation? Enter the slope as a fraction. -->(2/3)

OpenStudy (anonymous):

Manuel builds and sells wooden crates. He made a table to show how much his crates cost. Which equation represents the situation? Manuel’s Crates Crates 2 4 6 8 Cost 25 45 65 85 A. y – 2 = −10(x – 25) B. y – 4 = 10(x – 45) C. y – 65 = −10(x – 6) --> D. y – 85 = 10(x – 8)

OpenStudy (anonymous):

this is the question I got wrong it was d not c XD this one posted above this XD

OpenStudy (anonymous):

Carrington is paid at the same rate each time he baby sits. The graph shows the rates he is paid for different times. Which equation represents the situation? https://static.k12.com/calms_media/media/1512000_1512500/1512350/1/daf3a596b3d6b143c7a470822c2b47d5f1ee856f/MS_IMC-140516-100014.jpg the answer is y=7x+5

OpenStudy (anonymous):

Juan always saves the same amount from his weekly allowance. The table shows how much he has saved at different times. How much will Juan have saved in 14 wk? Juan’s Savings Weeks 3 6 9 12 Amount saved 19 28 37 46 this one was indeed $(52 )

OpenStudy (anonymous):

thanks for the help @AloneS :D

alones (alones):

what a work :)

OpenStudy (anonymous):

indeed it is... :D a looooot of hard math :( buuut, im done for today!

alones (alones):

lucky i'm still stuck on english

OpenStudy (anonymous):

need help?

OpenStudy (anonymous):

I also am stuck on a literature unit test... :(